OpenStudy (twistrose):

Dustin is mixing concrete. A formula for concrete calls for 3/8 gal of water. Dustin wants to make 7/9 more concrete than the formula creates. How much water should Dustin use? Express your answer as a fraction in simplest form.

OpenStudy (anonymous):

Multiply. \(\sf \dfrac{3}{8} \times \dfrac{7}{9}\)

OpenStudy (twistrose):

9*3=27 8*7=56 and you add it right

OpenStudy (twistrose):

@greenlegodude57

OpenStudy (anonymous):

No, you multiply the numerators together and then the denominators together.

OpenStudy (twistrose):

oh ok

OpenStudy (twistrose):

so 3*7=21 8*9=72

OpenStudy (anonymous):

Yes, that gives us: \(\sf\dfrac{21}{72}\)

OpenStudy (anonymous):

Now simplify it.

OpenStudy (twistrose):

oh ok thanks

OpenStudy (twistrose):

wait by what

OpenStudy (anonymous):

Find the factors of the 2 numbers and see if they have any in common.

OpenStudy (twistrose):

um ok can it be 3

OpenStudy (anonymous):

Yes, so divide them.

OpenStudy (anonymous):

21 / 3 = ? 72 / 3 = ?

OpenStudy (twistrose):

7/24